OpenStudy (anonymous):

what x-value makes this undefined f(x)=4x+1/x^2+5x+4

OpenStudy (solomonzelman):

Set the denominator equal to zero, and solve for x. \(\normalsize\color{blue}{ x^2+5x+4=0\LARGE\color{white}{ \rm │ }}\) (factor the left side)

OpenStudy (solomonzelman):

@CoolRanchDoritos , is there anything that is troubling you ?

OpenStudy (anonymous):

is it x+4 and x+1?

OpenStudy (solomonzelman):

When you factor, yes \(\normalsize\color{blue}{ x^2+5x+4=(x+4)(x+1)\LARGE\color{white}{ \rm │ }}\), so x=?

OpenStudy (solomonzelman):

what are the 2 solutions ?

OpenStudy (solomonzelman):

The 2 solutions you are going to get make the entire fraction undefined, because with them the denominator is equal to zero, and `anything / zero` = undefined.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{(x+4)(x+1)=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{x+4=0~~~~~~~~~or~~~~~~~~x+1=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{x=-4~~~~~~~~~~~~~~~~~~~~~~~~x=-1\LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

o ok thank you vary much